OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3

OpenStudy (akashdeepdeb):

Use the point-slope form a line. y-y1 = m(x-x1) You are given (x1,y1) = (8,-8) m = 3 And after doing that write it in the form of y=mx+b :D Understood? :)

OpenStudy (anonymous):

\[ y - y_1 = m \left( x - x_1\right) \text{, with point} \left( x_1 \text{, } y_1 \right)\]

OpenStudy (anonymous):

y-y1 = m(x-x1) Get the numbers in and answer !

OpenStudy (anonymous):

Okay yes I get it , it took me a while LOL

OpenStudy (anonymous):

and slope m. that is point-slope form

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

My final answer is y+8=3(x-8)

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Yes that is right. But they are asking us for it to be written in the slope-intercept form. Open the brackets and write it like this. y=mx+b :D

OpenStudy (anonymous):

8 = 3x+8 ?

OpenStudy (akashdeepdeb):

y+8=3(x-8) y+8 = 3x - 24 y = 3x - 32 <---- [This is the form in y=mx+b] Understood? :)

OpenStudy (anonymous):

Ah thanks okay yes I get it